# The geometry and arithmetic of bielliptic Picard curves

### v2

@@ 1.7. Acknowledgements
We thank Adam Morgan, Ciaran Schembri, Michael Stoll, and John Voight for helpful conversations and remarks. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157). Part of this research was carried out while the first author was visiting the second author in Jerusalem. We thank the Hebrew University of Jerusalem and the Einstein Institute of Mathematics for their hospitality.
@@ 3.    $(c)$
Remark 2.3.
A Picard curve is a plane quartic curve of the form (2.2). Such a curve is bielliptic (admitting a $\mu_{3}$-equivariant double cover to an elliptic curve) precisely when $c=0$, explaining our terminology.
@@ 2.5. The Prym variety $P$
This is an abelian surface. We refer to [27, §3.3] for more details concerning Prym varieties of bielliptic genus $3$ curves. The properties we will use are summarized in the following commutative diagram:
@@
where $d=a^{2}-4s^{3}$; this follows from [39, Theorem 1.1].
@@ 3.1. Generalities on $(1,2)$-polarized surfaces
This marks an interesting contrast with principal polarizations on abelian surfaces, which may not be represented by a $k$-rational line bundle. Indeed, this is the case for a positive proportion of genus 2 Jacobians over ${\rm\mathbb{Q}}$ ordered by height [37, Theorem 23].
@@ 3.2. Compatible $\mu_{3}$-actions on $(1,2)$-polarized abelian surfaces
To prove the claim about signatures, it suffices to show that the $\mu_{3}$-action on $T_{0}X_{1}$ and $T_{0}X_{2}$ act via inverse characters. Since $\mathscr{L}$ has self-intersection $4$ and $|\mathscr{L}|$ has four base points (Lemma 3.1), $X_{1}$ and $X_{2}$ intersect transversally at all these base points. Since $0\in A(k)$ is such a base point,
@@ 3.3. A Torelli theorem for marked bielliptic Picard curves
[…]
@@
Barth has shown that Torelli fails for the family of all bielliptic genus $3$ curves $(X,\tau)$. Suppose that $k$ is algebraically closed of characteristic zero and $A$ is an indecomposable $(1,2)$-polarized abelian surface $(A,\lambda)/k$. Then there exist many bielliptic genus $3$ curves $(X,\tau)$ over $k$ with dual Prym variety isomorphic to $(A,\lambda)$; by Proposition 3.4 any smooth member of the pencil $|\mathscr{L}|$ on $A$ will do. Theorem 3.13 shows that we can correct this when $C$ is a bielliptic Picard curve and we fix the $\mu_{6}$-action on $C$ and the $\mu_{3}$-action on $A$.
@@ 4.2. Moduli of marked $\mu_{3}$-abelian surfaces
The open substack $\mathfrak{Y}^{\text{indec}}\subset\mathfrak{Y}$ parametrizing those triples $(A,\lambda,\alpha)$ for which $(A,\lambda)$ is indecomposable has coarse space an open subset $Y^{\text{indec}}\subset Y$. Every decomposable $\mu_{3}$-abelian surface $(A,\lambda,\alpha)$ over ${\rm\mathbb{C}}$ is isomorphic to $(E\times E,\lambda_{O}\times 2\lambda_{O},\alpha_{1}\times\alpha_{2})$, where $E:y^{2}=x^{3}+1$, $\lambda_{O}$ is the unique principal polarization on $E$ and $\alpha_{i}$ are mutually inverse nontrivial $\mu_{3}$-actions on $E$. Since there are two choices for the pair $(\alpha_{1},\alpha_{2})$, even up to isomorphism, $(Y\setminus Y^{\text{indec}})({\rm\mathbb{C}})$ has size $2$. Since $Y$ is purely one-dimensional, $Y^{\text{indec}}$ is dense in $Y$. Theorem 3.13 and Lemma 2.8 show that the morphism ${\rm\mathbb{A}}^{2}_{{\rm\mathbb{Q}}}\setminus\{\Delta_{a,b}=0\}\rightarrow Y^{\text{indec}},(a,b)\mapsto[\operatorname{Prym}(C_{a,b},\gamma_{a,b})^{\vee}]$ is surjective, factors through the $j$-invariant morphism ${\rm\mathbb{A}}^{2}_{{\rm\mathbb{Q}}}\setminus\{\Delta_{a,b}=0\}\rightarrow{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}\setminus\{0,\infty\},(a,b)\mapsto j_{a,b}$ of (2.3), and induces an isomorphism ${\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}\setminus\{0,\infty\}\simeq Y^{\text{indec}},j\mapsto[\operatorname{Prym}(C_{a,b},\gamma_{a,b})^{\vee}]$, where $a,b$ are any elements with $j=j_{a,b}$. Putting everything together, $Y$ is a purely one-dimensional smooth variety having an open subset isomorphic to ${\rm\mathbb{P}}^{1}\setminus\{0,\infty\}$ whose complement has size $2$. We conclude that $Y\simeq{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}$ and that the isomorphism $j\colon Y^{\text{indec}}\rightarrow{\rm\mathbb{P}}_{{\rm\mathbb{Q}}}^{1}\setminus\{0,\infty\}$ uniquely extends to an isomorphism $Y\rightarrow{\rm\mathbb{P}}_{{\rm\mathbb{Q}}}^{1}$. ∎
@@ 4.3. Comparison with a quaternionic Shimura curve
We now compare $Y$ to a quaternionic Shimura curve, in the same spirit as [22, §4.2]. However, our analysis is a little different since we must consider twists and non-principal polarizations. We refer the reader to [45, Chapter 43] for background on Shimura curves.
@@
The nice curve $X$ is the coarse space of the moduli stack of triples $(A,\lambda,\iota)$, where $(A,\lambda)$ is a $(1,2)$-polarized abelian surface and $\iota\colon\mathcal{O}\rightarrow{\rm End}(A)$ is an embedding such that the Rosati involution restricts to $b\mapsto i\bar{b}i^{-1}$, where we recall from §4.1 that $i=1+2\omega$. (The moduli problem is typically formulated for principal polarizations, but the results of [45, §43.6] have analogues for the polarized order $(\mathcal{O},i)$, see [9, §12].) To such a triple $(A,\lambda,\iota)$ we would like to associate a $\mu_{3}$-abelian surface by restricting $\iota$ to ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\simeq\langle\omega\rangle\subset{\rm\mathbb{Z}}[\omega]^{\times}\subset\mathcal{O}^{\times}$. However, this defines a ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$-action on $A$; to get a $\mu_{3}$-action we need to suitably twist the situation.
@@
Corollary 4.8 was proved by Petkova and Shiga over ${\rm\mathbb{C}}$ using period matrix calculations [36, Proposition 7.1].
@@ 5. Identifying bielliptic Picard curves in the pencil
[…]
@@
[…]
@@ 6.1. Explicit quaternionic multiplication
To show that this automorphism is the identity, we will use that its construction can be carried out in families. More precisely, let $R={\rm\mathbb{Z}}[1/6,a,b,\Delta_{a,b}^{-1}]$ and $\tilde{R}=R[\varepsilon]/(\varepsilon^{6}-\Delta_{a,b})$. Equation (2.1) defines a universal marked bielliptic Picard curve $(\mathcal{C},\gamma)$ over $S={\rm Spec}(R)\subset{\rm\mathbb{A}}^{2}_{{\rm\mathbb{Z}}[1/6]}$ with Prym variety $(\mathcal{P},\lambda,\alpha)$ and dual $(\mathcal{A},\widehat{\lambda},\hat{\alpha})$. On the cover $\tilde{S}={\rm Spec}(\tilde{R})\rightarrow S$ given by adjoining a sixth root of $\Delta_{a,b}$, there exists an isomorphism $\Psi\colon\mathcal{A}_{\tilde{S}}\rightarrow\mathcal{P}_{\tilde{S}}$ with the property that $\Psi$ specializes to $\psi$ for all morphisms ${\rm Spec}(k)\rightarrow\tilde{S}$ and that $(\Psi^{\vee})^{-1}\circ\Psi\in{\rm Aut}_{\tilde{S}}(\mathcal{A},\lambda,\alpha)$. Since the moduli stack of $(1,2)$-polarized abelian surfaces is separated and Deligne–Mumford, $\mathbf{Aut}_{\tilde{S}}(\mathcal{A},\lambda,\alpha)\rightarrow\tilde{S}$ is finite and unramified. Since finite unramified morphisms are étale locally on the target disjoint unions of closed immersions [44, Tag 04HJ] and the base $\tilde{S}$ is connected, it suffices to show that $(\Psi^{\vee})^{-1}\Psi$ is the identity for a single specialization $s\in\tilde{S}$.
@@
Let $C/k$ be a bielliptic Picard curve with Prym variety $P$ and suppose that $k$ is algebraically closed. Then $P$ is not simple normal-$\Leftrightarrow$ $\mathcal{O}\not\simeq{\rm End}(P)$ normal-$\Leftrightarrow$ $P$ is isogenous to the square of a CM elliptic curve. Consequently, if $P$ is simple then ${\rm End}(P)=\iota_{x}(\mathcal{O})$ for every $x\in\mathcal{S}$.
@@
If $\mathrm{char}(k)=0$ or if $k=\overline{{\rm\mathbb{F}}}_{p}$ with $p>3$, then these conditions are also equivalent to the condition that $P$ is isomorphic to a product of elliptic curves. See [40] for the former case and [38, Theorem 3.1] for the latter.
@@ 6.4. Geometrically split Prym varieties
In the coordinates $j$ and $t$ of $Y$ and $X^{*}$ respectively, the map $\Pi\colon Y\rightarrow X^{*}$ is given by $j\mapsto\frac{(j+1)^{2}}{4j}$. A $k$-point $t\in k$ on $X^{*}$ lifts to $Y$ if and only if $t(t-1)$ is a square in $k$.
@@ 7.2. $\sqrt{-3}$-torsion
We use the $\mu_{3}$-action on $P$ to explicitly analyze a ${\rm Gal}_{k}$-submodule of $P[3]$. Let $\mathfrak{p}$ be the ideal $(1-\omega)\subset{\rm\mathbb{Z}}[\omega]$, viewed as a subgroup of ${\rm End}(P_{k^{\operatorname{sep}}})$. Since $\mathfrak{p}$ is ${\rm Gal}_{k}$-stable, so is the subgroup $A[\mathfrak{p}]$ of points fixed by $\omega$. We have $\mathfrak{p}=(\sqrt{-3})$, so $P[\mathfrak{p}]$ has order $9$, and is the kernel of an isogeny $P\to B$ over $k$. In fact, the quotient $B=P/P[\mathfrak{p}]$ is isomorphic to the sextic twist $P_{-27}$, which is also the quadratic twist of $P$ by the extension $k(\omega)/k$ [41, Rem 2.8]. We therefore have an exact sequence

### v3

@@ 1.7. Acknowledgements
We thank Adam Morgan, Ciaran Schembri, Michael Stoll, John Voight for helpful conversations and remarks. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157). Part of this research was carried out while the first author was visiting the second author in Jerusalem. We thank the Hebrew University of Jerusalem and the Einstein Institute of Mathematics for their hospitality.
@@ 3.    $(c)$
We say $C$ a nice genus $3$ curve is a Picard curve if it admits a faithful $\mu_{3}$-action such that the quotient $C/\mu_{3}$ is isomorphic to ${\rm\mathbb{P}}^{1}$. We say $C$ is a bielliptic curve if it admits an involution $\tau$ such that the quotient $C/\tau$ is a genus $1$ curve. To avoid notational conflicts, we show:
Lemma 2.4.
If a nice genus $3$ curve $C/k$ is both a Picard curve and a bielliptic curve, then it is a bielliptic Picard curve in the sense of Definition 2.1.
Proof.
Since $C$ is a Picard curve, it has an affine equation of the form $y^{3}=f(x)$ for some monic quartic $f(x)=x^{4}+ax^{2}+bx+c\in k[x]$ with nonzero discriminant, and the $\mu_{3}$-action on $C$ is given by $\omega\cdot(x,y)=(x,\omega y)$. By assumption, there exists an involution $\tau\colon C\rightarrow C$ such that $C/\tau$ has genus $1$. Let $C_{0}$ be the Picard curve with equation $y^{3}=x^{4}+1$. If $C_{k^{\operatorname{sep}}}$ is not isomorphic to $C_{0,k^{\operatorname{sep}}}$, then a calculation using [10, Lemma 1.21(a)] implies that $b=0$ and that $\tau$ maps $(x,y)$ to $(-x,y)$, so $C$ is indeed a bielliptic Picard curve. We may therefore assume that $C_{k^{\operatorname{sep}}}\simeq C_{0,k^{\operatorname{sep}}}$. Since every two order $3$ subgroups of ${\rm Aut}(C_{0,k^{\operatorname{sep}}})$ are conjugate (by an explicit computation in the group (48,33)), there is an isomorphism $C_{k^{\operatorname{sep}}}\xrightarrow{\sim}C_{0,k^{\operatorname{sep}}}$ that preserves the $\mu_{3}$-actions on both sides. Such an isomorphism must be of the form $(x,y)\mapsto(\alpha x,\beta y)$ for some $\alpha,\beta\in k^{\operatorname{sep}}\setminus\{0\}$, using the same reasoning as the proof of [10, Lemma 1.21]. It follows that $a=b=0$ and $C:y^{3}=x^{4}+c$ is a (special) bielliptic Picard curve. ∎
@@ 2.5. The Prym variety $P$
This is an abelian surface. We refer to [27, §3.3] for more details concerning Prym varieties of bielliptic genus $3$ curves. The properties we will use can be deduced from [35] and are summarized in the following commutative diagram:
@@
where $d=a^{2}-4s^{3}$; this follows from [40, Theorem 1.1].
@@ 3.1. Generalities on $(1,2)$-polarized surfaces
This marks an interesting contrast with principal polarizations on abelian surfaces, which may not be represented by a $k$-rational line bundle. Indeed, this is the case for a positive proportion of genus 2 Jacobians over ${\rm\mathbb{Q}}$ ordered by height [38, Theorem 23].
@@ 3.2. Compatible $\mu_{3}$-actions on $(1,2)$-polarized abelian surfaces
To prove the claim about signatures, it suffices to show that the $\mu_{3}$-action on $T_{0}X_{1}$ and $T_{0}X_{2}$ acts via inverse characters. Since $\mathscr{L}$ has self-intersection $4$ and $|\mathscr{L}|$ has four base points (Lemma 3.1), $X_{1}$ and $X_{2}$ intersect transversally at all these base points. Since $0\in A(k)$ is such a base point,
@@ 3.3. A Torelli theorem for marked bielliptic Picard curves
(3.3) $\begin{Bmatrix}\text{Marked bielliptic}\\ \text{Picard curves over }k\end{Bmatrix}\rightarrow\begin{Bmatrix}\text{Indecomposable}\\ \mu_{3}\text{-abelian surfaces over }k\end{Bmatrix}.$
@@
Barth has shown that a Torelli-type result fails for the family of all bielliptic genus $3$ curves $(X,\tau)$. Suppose that $k$ is algebraically closed of characteristic zero and $A$ is an indecomposable $(1,2)$-polarized abelian surface $(A,\lambda)/k$. Then there exist many bielliptic genus $3$ curves $(X,\tau)$ over $k$ with dual Prym variety isomorphic to $(A,\lambda)$; by Proposition 3.4 any smooth member of the pencil $|\mathscr{L}|$ on $A$ will do. Theorem 3.13 shows that we can correct this when $C$ is a bielliptic Picard curve and we fix the $\mu_{6}$-action on $C$ and the $\mu_{3}$-action on $A$.
@@ 4.2. Moduli of marked $\mu_{3}$-abelian surfaces
The open substack $\mathfrak{Y}^{\text{indec}}\subset\mathfrak{Y}$ parametrizing those triples $(A,\lambda,\alpha)$ for which $(A,\lambda)$ is indecomposable has coarse space an open subset $Y^{\text{indec}}\subset Y$. We first analyze the complement. Every decomposable $(1,2)$-polarized abelian surface $(A,\lambda)$ over $\mathbb{C}$ must be of the form $(E_{1},\lambda_{O})\times(E_{2},2\lambda_{O})$, where $E_{i}$ are elliptic curves and $\lambda_{O}$ is the unique principal polarization on an elliptic curve. If $\alpha$ is a $\mu_{3}$-action on $A$ satisfying the conditions of Definition 3.7, then $\alpha$ restricts to a nontrivial action on $E_{i}$ for each $i$. It follows that $E_{i}$ must be isomorphic to $E\colon y^{2}=x^{3}+1$. We conclude that every decomposable $\mu_{3}$-abelian surface $(A,\lambda,\alpha)$ over ${\rm\mathbb{C}}$ is isomorphic to $(E\times E,\lambda_{O}\times 2\lambda_{O},\alpha_{1}\times\alpha_{2})$, where $\alpha_{i}$ are mutually inverse nontrivial $\mu_{3}$-actions on $E$. Since there are two choices for the pair $(\alpha_{1},\alpha_{2})$, even up to isomorphism, $(Y\setminus Y^{\text{indec}})({\rm\mathbb{C}})$ has size $2$. Since $Y$ is purely one-dimensional, $Y^{\text{indec}}$ is dense in $Y$. Theorem 3.13 and Lemma 2.8 show that the morphism ${\rm\mathbb{A}}^{2}_{{\rm\mathbb{Q}}}\setminus\{\Delta_{a,b}=0\}\rightarrow Y^{\text{indec}},(a,b)\mapsto[\operatorname{Prym}(C_{a,b},\gamma_{a,b})^{\vee}]$ is surjective, factors through the $j$-invariant morphism ${\rm\mathbb{A}}^{2}_{{\rm\mathbb{Q}}}\setminus\{\Delta_{a,b}=0\}\rightarrow{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}\setminus\{0,\infty\},(a,b)\mapsto j_{a,b}$ of (2.3), and induces an isomorphism ${\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}\setminus\{0,\infty\}\simeq Y^{\text{indec}},j\mapsto[\operatorname{Prym}(C_{a,b},\gamma_{a,b})^{\vee}]$, where $a,b$ are any elements with $j=j_{a,b}$. Putting everything together, $Y$ is a purely one-dimensional smooth variety having an open subset isomorphic to ${\rm\mathbb{P}}^{1}\setminus\{0,\infty\}$ whose complement has size $2$. We conclude that $Y\simeq{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}$ and that the isomorphism $j\colon Y^{\text{indec}}\rightarrow{\rm\mathbb{P}}_{{\rm\mathbb{Q}}}^{1}\setminus\{0,\infty\}$ uniquely extends to an isomorphism $Y\rightarrow{\rm\mathbb{P}}_{{\rm\mathbb{Q}}}^{1}$. ∎
@@ 4.3. Comparison with a quaternionic Shimura curve
We now compare $Y$ to a quaternionic Shimura curve, in the same spirit as [22, §4.2]. However, our analysis is a little different since we must consider twists and non-principal polarizations. We refer the reader to [46, Chapter 43] for background on Shimura curves.
@@
The nice curve $X$ is the coarse space of the moduli stack of triples $(A,\lambda,\iota)$, where $(A,\lambda)$ is a $(1,2)$-polarized abelian surface and $\iota\colon\mathcal{O}\rightarrow{\rm End}(A)$ is an embedding such that the Rosati involution restricts to $b\mapsto i\bar{b}i^{-1}$, where we recall from §4.1 that $i=1+2\omega$. (The moduli problem is typically formulated for principal polarizations, but the results of [46, §43.6] have analogues for the polarized order $(\mathcal{O},i)$, see [9, §12].) To such a triple $(A,\lambda,\iota)$ we would like to associate a $\mu_{3}$-abelian surface by restricting $\iota$ to ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\simeq\langle\omega\rangle\subset{\rm\mathbb{Z}}[\omega]^{\times}\subset\mathcal{O}^{\times}$. However, this defines a ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$-action on $A$; to get a $\mu_{3}$-action we need to suitably twist the situation.
@@
Corollary 4.8 was proved by Petkova and Shiga over ${\rm\mathbb{C}}$ using period matrix calculations [37, Proposition 7.1].
@@ 5. Identifying bielliptic Picard curves in the pencil
Suppose that $K\neq k$ and let $\sigma\in{\rm Gal}(K/k)$ be the unique nontrivial element. Let $\chi=\psi|_{F_{2p}}\circ\phi_{p}\colon C_{\text{twist},K}\xrightarrow{\sim}\mathcal{Q}_{K}$. Then $\chi^{\sigma}\circ\chi^{-1}\colon\mathcal{Q}_{K}\rightarrow\mathcal{Q}_{K}$ equals the involution $(u,v)\mapsto(-u,v)$.
@@
Let $C$ be a bielliptic Picard curve of the form $C_{a,b}$ for some $a,b\in k$. Then $C_{\text{twist}}$ is again a bielliptic Picard curve, isomorphic to $C_{\delta a,\delta^{2}b}$ where $\delta=\Delta_{a,b}=16b(a^{2}-4b)$. In particular, $C$ and $C_{\text{twist}}$ are isomorphic over $k^{\operatorname{sep}}$.
@@ 6.1. Explicit quaternionic multiplication
To show that this automorphism is the identity, we will use that its construction can be carried out in families. More precisely, let $R={\rm\mathbb{Z}}[1/6,a,b,\Delta_{a,b}^{-1}]$ and $\tilde{R}=R[\varepsilon]/(\varepsilon^{6}-\Delta_{a,b})$. Equation (2.1) defines a universal marked bielliptic Picard curve $(\mathcal{C},\gamma)$ over $S={\rm Spec}(R)\subset{\rm\mathbb{A}}^{2}_{{\rm\mathbb{Z}}[1/6]}$ with Prym variety $(\mathcal{P},\lambda,\alpha)$ and dual $(\mathcal{A},\widehat{\lambda},\hat{\alpha})$. On the cover $\tilde{S}={\rm Spec}(\tilde{R})\rightarrow S$ given by adjoining a sixth root of $\Delta_{a,b}$, there exists an isomorphism $\Psi\colon\mathcal{A}_{\tilde{S}}\rightarrow\mathcal{P}_{\tilde{S}}$ with the property that $\Psi$ specializes to $\psi$ for all morphisms ${\rm Spec}(k)\rightarrow\tilde{S}$ and that $(\Psi^{\vee})^{-1}\circ\Psi\in{\rm Aut}_{\tilde{S}}(\mathcal{A},\lambda,\alpha)$. Since the moduli stack of $(1,2)$-polarized abelian surfaces is separated and Deligne–Mumford, $\mathbf{Aut}_{\tilde{S}}(\mathcal{A},\lambda,\alpha)\rightarrow\tilde{S}$ is finite and unramified. Since finite unramified morphisms are étale locally on the target disjoint unions of closed immersions [45, Tag 04HJ] and the base $\tilde{S}$ is connected, it suffices to show that $(\Psi^{\vee})^{-1}\Psi$ is the identity for a single specialization $s\in\tilde{S}$.
@@
Let $C/k$ be a bielliptic Picard curve with Prym variety $P$ and suppose that $k$ is algebraically closed. Then $P$ is not simple $\Leftrightarrow$ $\mathcal{O}\not\simeq{\rm End}(P)$Leftrightarrow $P$ is isogenous to the square of a CM elliptic curve. Consequently, if $P$ is simple then ${\rm End}(P)=\iota_{x}(\mathcal{O})$ for every $x\in\mathcal{S}$.
@@
If $\mathrm{char}(k)=0$ or if $k=\overline{{\rm\mathbb{F}}}_{p}$ with $p>3$, then these conditions are also equivalent to the condition that $P$ is isomorphic to a product of elliptic curves. See [41] for the former case and [39, Theorem 3.1] for the latter.
@@ 6.4. Geometrically split Prym varieties
In the coordinates $j$ and $t$ of $Y$ and $X^{*}$ respectively, the map $\Pi\colon Y\rightarrow X^{*}$ is given by $j\mapsto\frac{(j+1)^{2}}{4j}$. A $k$-point $t\in k$ on $X^{*}$ lifts to a $k$-point on $Y$ if and only if $t(t-1)$ is a square in $k$.
@@ 7.2. $\sqrt{-3}$-torsion
We use the $\mu_{3}$-action on $P$ to explicitly analyze a ${\rm Gal}_{k}$-submodule of $P[3]$. Let $\mathfrak{p}$ be the ideal $(1-\omega)\subset{\rm\mathbb{Z}}[\omega]$, viewed as a subgroup of ${\rm End}(P_{k^{\operatorname{sep}}})$. Since $\mathfrak{p}$ is ${\rm Gal}_{k}$-stable, so is the subgroup $A[\mathfrak{p}]$ of points fixed by $\omega$. We have $\mathfrak{p}=(\sqrt{-3})$, so $P[\mathfrak{p}]$ has order $9$, and is the kernel of an isogeny $P\to B$ over $k$. In fact, the quotient $B=P/P[\mathfrak{p}]$ is isomorphic to the sextic twist $P_{-27}$, which is also the quadratic twist of $P$ by the extension $k(\omega)/k$ [42, Rem 2.8]. We therefore have an exact sequence